# Entropy-Regularized Process Reward Model

### v1

@@ 5.2 RLHF Results
Table 1 summarizes the performance improvements using RLHF enhanced via ER-PRM and baseline PRMs. Compared to the baselines, Our ER-PRM outperformed by surpassing Soft-Label PRM 0.8% and 2.7% on GSM8K and MATH, and Hard-Label PRM 0.5% and 0.8%, respectively for the Mistral model. We also observe a larger improvement for the deepseek model with respect to the baselines, indicating ER-PRM’s superior capability in guiding the optimal policy model and its robustness in handling more complex task settings. We also notice that the Hard-Label setting consistently outperforms the Soft-Label setting in RLHF. The surprisingly strong performance for Hard-Label indicates its feasibility, also a different discovery compared to Luo et al. (2024).
@@
In general, our RLHF approach achieved superior performance in evaluating the policy model in GSM8K and MATH. The reason is that our ER-PRM consistently having 1-2% higher best-of-N accuracy in selecting data from both in-domain and out-of-domain datasets than the baselines, showcasing the stability and effectiveness of our entropy-regularized reward formulation in selecting accurate rationales from varying quality data sources.
@@ 6 Analysis
6.1 Entropy-Regularized Reward Analysis
@@
The result is shown in Figure LABEL:fig:mistral-policy-best-of-n-eta for best-of-N performance on each $\eta$. We observe that there is a peak performance for each dataset and then the performance drops when $\eta$ is too large or small. The experiments show that $\eta=1$ or $\eta=2$ is the best for hard dataset MATH. $\eta=15$ is the best for easy dataset GSM8K. Larger $\eta$ represents a larger penalty on the entropy for the process rewards where they are pushed higher and closer to the hard label estimation, while smaller $\eta$ represents a small penalty on the entropy for the process rewards. We conclude that to achieve the best performance on automatic process reward labeling, a more fine-grained estimation of the reward scores should be applied. Simply applying hard label estimation may introduce too much noise for reward model training. We also notice that vanilla soft label estimation may not be an optimal strategy, as the performance is relatively bad when $\eta$ is small ($\eta=0.1$) for both datasets.
@@ 6.2 Regression Training Strategy
In addition to training the reward model in an auto-regressive way, we also explore the reward models with the regression training strategy, where we train the model to output a score directly given a problem and current steps. The best-of-N evaluation results are shown in Figure LABEL:fig:regression. The performance of best-of-N increases as N becomes larger. Our reward model still outperforms the Soft-label and Hard-label settings. We also identify it as a method to train reward models. However, we notice that auto-regressive training could yield better results for reward models than regression training. On both GSM8K and MATH datasets, the best-of-N accuracies are consistently lower than the models trained using the auto-regressive strategy. Especially on MATH, there is a significant gap of the accuracy between the two training methods. One hypothesis is that auto-regressive training integrates the steps, reducing the total number of training steps, which in turn leads to less forgetting. For the MATH dataset, since there are more reasoning steps per question, splitting them into multiple training samples increases the difficulty for the reward models to learn. It is also possible that the reasoning steps in auto-regressive training are continuous and coherent, allowing the reward model to better learn the process of assigning rewards.
@@ 7 Conclusion
8 Limitations
Although our method demonstrates the potential to improve the reward models and the policy models with rejection fine-tuning, we have not conducted large-scale online RLHF experiments such as PPO (Schulman et al., 2017), due to the limitation of the computational resources. Future work could explore more policy models training for mathematical reasoning using more RLHF methods, combined with ER-PRM.
Appendix A Technical Lemma
@@
Appendix B Hyper-parameter Settings
We train the reward models using meta-llama/Llama-3.1-8B Dubey et al. (2024) as the base model. The reward models are trained with a global batch size of 64, a learning rate of 1e-6, a max length of 512, and an epoch of 1 on 4 H100 GPUs. We have conducted hyper-parameter searching on the learning rate on {$5e^{-6}$,$2e^{-6}$,$1e^{-6}$,$5e^{-7}$}. During the automatic labeling process, we use an $\eta$ of 2 in our main experiments for Mitral data, and an $\eta$ of 10 for deepseek data. We also conducted hyper-parameter searching on the $\eta$ on {$0.1$, $1$, $2$, $5$, $8$, $10$, $15$}.
@@ Appendix C Best-of-N Evaluation Results in Tables
In this section, we list the performance of the best-of-N evaluation of ER-PRM and other baselines in Tables. The results for the reward models with policy model Mistral-MetaMath-7b are shown in Table 2, which corresponds to Figure LABEL:fig:mistral-policy-best-of-n. The results for the reward models with policy model deepseek-math-7b-instruct are shown in Table 3, which corresponds to Figure LABEL:fig:deepseek-instruct-policy-best-of-n.
@@
Appendix D Reward Models with Various Sizes
In addition to training the reward model with Llama-3.1-8B, we also explore training on various sizes. In this section, we train the reward model with Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct (Dubey et al., 2024). We hope to increase the performance of smaller-size models so we use the instruct-version for them instead of base models. We train them using the Mistral data, the same as the training data used in the main experiments.
The experiment results are shown in Table 7 and Table 8. Although we change into models of smaller sizes, the best-of-N accuracy still outperforms the Top 1 accuracy by a large margin. For example, for the 1B size model, the best-of-4 accuracy outperforms Top 1 accuracy by 4% and the best-of-64 accuracy outperforms by 5$\sim$6on GSM8K dataset. We observe that the performance gain from best-of-4 to best-of-64 for the 1B model is much smaller than the 8B model shown in Table 2, which has 4$\sim$6performance gain while 1B model has only 2$\sim$3performance gain. The comparison demonstrates that larger models have higher chances of selecting the gold answer from candidates when $N$ is large. We also discover that the 1B model suffers from severe reward hacking problems on the MATH test set. For ER-PRM and Hard-label PRM, they both achieve an accuracy of 37.6% on best-of-16 evaluation. However, their accuracy drops about 2% on best-of-64 evaluation, showing that they are distracted by the wrong answers and could not effectively select the gold responses. The reward hacking problem becomes less severe for the easy dataset GSM8K, which probably due to the less number of wrong candidate answers.
For the 3B size model, it consistently outperforms the 1B size model, showing that training with larger-size reward models would generally achieve better performance. The 3B reward model on MATH shows a competent performance as the 8B reward model (Table 2), demonstrating the feasibility of training small-size but powerful models. We also discover that the reward hacking problem is largely mitigated for the 3B size model, as the best-of-N accuracy keeps increasing on MATH as $N$ increases, and does not drop much on GSM8K from best-of-32 to best-of-64. However, we still notice that on the GSM8K test set, the performance gain from $N=4$ to $N=64$ is still little compared to the 8B case. We may conclude that a larger model size reward model would generally have a larger performance gain when $N$ increases.
GSM8K MATH500
N=4 82.41 82.56 82.18 32.6 32.8 31.4
N=8 82.56 82.56 82.41 34.8 33.0 33.6
N=16 83.70 82.94 83.54 37.6 35.2 37.6
N=32 84.31 84.00 83.47 36.4 35.8 36.2
N=64 84.08 83.93 83.47 35.8 35.0 35.4
Top1@acc 78.09 29.4
GSM8K MATH500
N=4 83.54 83.32 83.39 36.0 36.0 35.4
N=8 83.92 83.47 83.47 37.8 37.2 37.2
N=16 84.23 84.00 83.92 38.4 38.2 37.4
N=32 84.91 84.68 84.38 41.8 39.2 40.2
N=64 85.36 85.06 84.68 41.8 39.6 41.0
Top1@acc 78.09 29.4
Appendix E OOD Performance of PRMs
In this section, we examine the OOD performance of the PRMs. We use deepseek-math-7b-rl (Shao et al., 2024) as the policy model for best-of-N generation. We train the reward model from Llama-3.1-8b on Mistral data.
The results are shown in Figure LABEL:fig:deepseek-policy-best-of-n and Table 9. The accuracy increases slowly when N increases from 4 to 64 on GSM8K. And the increment on the MATH dataset is also slower than in the case where we use Mistral as the policy model. We observe the performance drop for ORM reward models, indicating its instability for out-of-distribution generated data. The reason for the smaller performance gain is that, our process reward data is generated from a weaker model Mistral-MetaMath-7b. The reward model might only understand the behavior of the weak model. For certain steps where the weak model fails to produce the correct answer, the reward model is likely to assign these steps a low score. However, a stronger model may be capable of deriving the correct answer in such cases.
@@ Appendix E OOD Performance of PRMs
Appendix F OpenMath2-Llama3.1-8B as the Policy Model on MATH
In this section, We want to examine the reward model performance paired with one of the most powerful open-source policy models fine-tuned from open-source data. We could also see the reward model performance on the Out-of-Distribution Scenario. We use OpenMath2-Llama3.1-8B (Toshniwal et al., 2024b) as the generator, which is fine-tuned from 14 million mathematical problems and excellent in mathematical reasoning. It achieves a Top 1 accuracy of 66.2% of the MATH500 test set. We evaluate the best-of-N performance on the MATH500 test dataset. For each question, we use OpenMath2-Llama3.1-8B to generate N candidate solutions and use the reward model to select the response. We use the reward model trained from Llama-3.1-8B on deepseek data.
MATH
Best-of-N Ours Soft-label PRM Hard-label PRM
N=4 69.0 67.2 69.2
N=8 70.6 68.6 68.8
N=16 70.4 69.2 69.4
N=32 70.2 68.4 69.8
N=64 70.4 67.6 68.6
Appendix G Experiments on Entire MATH Test Set
In our previous experiments, we evaluate the best-of-N performance on the MATH500 test sets to align with previous works (Wang et al., 2024; Lightman et al., 2023) and save computation cost. In this section, we evaluate our reward models on the entire MATH test set which contains 5000 questions. The reward models are trained from Llama-3.1-8B on Mistral data. And the Policy model for best-of-N generation is Mistral-MetaMath-7b.
The results are shown in Table 11. When evaluated on the MATH test set, we could still observe an increasing trend in the accuracy as N increases. Compared to Hard-label and Soft-label settings, our model consistently outperforms them. When N is larger, the increment for Soft-label and Hard-label PRM is less noticeable while the increment for ER-PRM is still significant.
MATH
Best-of-N Ours Soft-label PRM Hard-label PRM
N=4 32.94 32.80 33.38
N=8 34.72 33.80 34.28
N=16 36.24 35.00 36.18
N=32 37.80 36.12 37.46
N=64 38.62 36.74 37.50
Table 11: Best-of-N evaluation on the entire MATH test set. The reward models are trained from Llama-3.1-8B on Mistral data. The policy model for generation is Mistral-MetaMath-7b.
Appendix H Definition of Step, Partial Chain, and Chain

### v2

@@ 5.3 Performance Evaluation
Table 1 summarizes the performance improvements achieved using RLHF enhanced via ER-PRM compared to baseline PRMs. ER-PRM outperforms the baselines, surpassing Soft-Label PRM by 0.8% and 2.7% on GSM8K and MATH, respectively, and Hard-Label PRM by 0.5% and 0.8% for the Mistral model. We observe even larger improvements for the DeepSeek model compared to the baselines, highlighting ER-PRM’s superior ability to guide the more potent policy model and its robustness in handling more complex task settings. Additionally, we find that the Hard-Label setting consistently outperforms the Soft-Label setting in RLHF, with the surprisingly strong performance of Hard-Label indicating its feasibility. This is a notable discovery, differing from the findings of Luo et al. (2024).
@@
In general, our RLHF approach achieved superior performance in evaluating the policy model on GSM8K and MATH test sets. This improvement can be attributed to our ER-PRM, which consistently demonstrates 1-2% higher best-of-N accuracy in selecting data from both in-domain and out-of-domain datasets compared to the baselines. This highlights the stability and effectiveness of our entropy-regularized reward formulation in selecting accurate rationales from varying quality data sources.
@@ 6 Analysis
6.1 Reward Models with Various Sizes
In addition to training the reward model with Llama-3.1-8B, we also explore training on various sizes. In this section, we train the reward model with Llama-3.2-3B-Instruct and Llama-3.2-1B-Instruct (Dubey et al., 2024). We hope to increase the performance of smaller-size models so we use the instruct-version for them instead of base models. We train both of them using the Mistral data.
Experiments with Llama-3.2-1B-Instruct
[…]
Experiments with Llama-3.2-3B-Instruct
The experiment results are shown in Table 3. For the 3B size model, it consistently outperforms the 1B size model, showing that training with larger-size reward models would generally achieve better performance. The 3B reward model on MATH shows a competent performance as the 8B reward model (Table 5), demonstrating the feasibility of training small-size but powerful models. We also discover that the reward hacking problem is largely mitigated for the 3B size model, as the best-of-N accuracy keeps increasing on MATH as $N$ increases, and does not drop much on GSM8K from best-of-32 to best-of-64. However, we still notice that on the GSM8K test set, the performance gain from $N=4$ to $N=64$ is still little compared to the 8B case. We may conclude that a larger model size reward model would generally have a larger performance gain when $N$ increases.
GSM8K MATH500
N=4 82.1±0.2 82.2±0.3 81.9±0.2 32.3±0.5 32.0±0.6 31.5±0.5
N=8 82.6±0.1 82.3±0.2 82.2±0.4 34.2±0.5 33.7±0.5 33.5±0.4
N=16 83.5±0.3 83.2±0.2 82.8±0.3 37.0±0.6 35.4±0.6 36.3±0.7
N=32 84.1±0.3 83.6±0.5 83.8±0.4 36.4±0.8 35.9±0.7 36.2±1.1
N=64 84.0±0.3 83.8±0.4 83.9±0.4 36.3±0.9 35.5±0.9 35.9±1.3
Top1@acc 78.09 29.4
GSM8K MATH500
N=4 83.6±0.3 83.3±0.3 83.3±0.3 35.6±0.6 35.2±0.4 35.3±0.7
N=8 84.1±0.4 83.7±0.5 83.5±0.4 37.7±0.8 37.1±0.8 37.4±0.5
N=16 84.5±0.3 84.0±0.2 83.9±0.4 38.9±0.6 38.2±0.7 38.3±0.6
N=32 85.0±0.2 84.6±0.4 84.4±0.5 40.8±1.1 39.6±0.9 40.1±0.8
N=64 85.4±0.4 85.0±0.4 84.8±0.3 41.9±1.0 40.2±0.8 41.1±0.8
Top1@acc 78.09 29.4
6.2 Out-of-Distribution Performance of Process Reward Models
In this subsection, we examine the OOD performance of the PRMs. The DeepSeek-Math-7B-RL (Shao et al., 2024) and OpenMath2-Llama3.1-8B (Toshniwal et al., 2024b) are used as the policy models for generation, and the reward models are used for best-of-N evaluation.
DeepSeek-Math-7B-RL as the policy model
We firstly use DeepSeek-Math-7B-RL as the policy model for best-of-N generation. We train the reward model from Llama-3.1-8b on Mistral data. The results are shown in Figure 5(a) and Table 10. The accuracy increases slowly when N increases from 4 to 64 on GSM8K. And the increment on the MATH dataset is also slower than in the case where we use Mistral as the policy model. We observe the performance drop for ORM reward models, indicating its instability for out-of-distribution generated data. The reason for the smaller performance gain is that, our process reward data is generated from a weaker model Mistral-MetaMath-7b. The reward model might only understand the behavior of the weak model. For certain steps where the weak model fails to produce the correct answer, the reward model is likely to assign these steps a low score. However, a stronger model may be capable of deriving the correct answer in such cases.
(a) Best-of-N evaluation results on GSM8K and MATH500 datasets with DeepSeek-Math-7B-RL as the generator. The Hard-label PRM is the same setting as Math-Shepherd.
OpenMath2-Llama3.1-8B as the policy model
We want to examine the reward model performance paired with one of the most powerful open-source policy models fine-tuned from open-source data. Especially, we hope to lift the performance on MATH, which is difficult for most of the open-source models. We use OpenMath2-Llama3.1-8B (Toshniwal et al., 2024b) as the generator, which is fine-tuned from 14 million mathematical problems and excellent in mathematical reasoning. It achieves a Top 1 accuracy of 66.2% of the MATH500 test set. We evaluate the best-of-N performance on the MATH500 test dataset. For each question, we use OpenMath2-Llama3.1-8B to generate N candidate solutions and use the reward model to select the response. We use the reward model trained from Llama-3.1-8B on DeepSeek-data.
6.3 Impact of Entropy Regularization
(b) Best-of-N evaluation results on GSM8K and MATH500 datasets with Mistral-MetaMath-7b as the generator with different $\eta$. The reward models are trained on Mistral data on GSM8K and MATH separately.
(c) Best-of-N evaluation results on GSM8K and MATH500 datasets with Mistral-MetaMath-7b as the generator. The reward models are trained using regression. They are also compared with the auto-regressive training strategy. The reward models are all trained on Mistral data.
@@ 6.3 Impact of Entropy Regularization
The result is shown in Figure 5(b) for best-of-N performance on each $\eta$. We observe that there is a peak performance for each dataset and then the performance drops when $\eta$ is too large or small. The experiments show that $\eta=1$ or $\eta=2$ is the best for hard dataset MATH. $\eta=15$ is the best for easy dataset GSM8K. Larger $\eta$ represents a larger penalty on the entropy for the process rewards where they are pushed higher and closer to the hard label estimation, while smaller $\eta$ represents a small penalty on the entropy for the process rewards. We conclude that to achieve the best performance on automatic process reward labeling, a more fine-grained estimation of the reward scores should be applied. Simply applying hard label estimation may introduce too much noise for reward model training. We also notice that vanilla soft label estimation may not be an optimal strategy, as the performance is relatively bad when $\eta$ is small ($\eta=0.1$) for both datasets.
@@ 6.4 Regression Training Strategy
In addition to training the reward model in an auto-regressive way, we also explore the reward models with the regression training strategy, where we train the model to output a score directly given a problem and current steps. The best-of-N evaluation results are shown in Figure 5(c).
MATH
Best-of-N Ours Soft-label PRM Hard-label PRM
N=4 69.1±0.4 67.5±0.6 69.2±0.3
N=8 70.4±0.6 69.1±0.6 69.2±0.5
N=16 70.5±0.5 69.3±0.8 69.7±0.6
N=32 70.3±0.4 68.8±0.7 69.6±0.7
N=64 70.4±0.6 68.3±0.5 68.8±0.4
The performance of best-of-N increases as N becomes larger. Our reward model still outperforms the Soft-label and Hard-label settings. We also identify it as a method to train reward models. However, we notice that auto-regressive training could yield better results for reward models than regression training. On both GSM8K and MATH datasets, the best-of-N accuracies are consistently lower than the models trained using the auto-regressive strategy. Especially on MATH, there is a significant gap of the accuracy between the two training methods. One hypothesis is that auto-regressive training integrates the steps, reducing the total number of training steps, which in turn leads to less forgetting. For the MATH dataset, since there are more reasoning steps per question, splitting them into multiple training samples increases the difficulty for the reward models to learn. It is also possible that the reasoning steps in auto-regressive training are continuous and coherent, allowing the reward model to better learn the process of assigning rewards.
@@ 7 Conclusion
Appendix A Ethical Statements and Limitations
Our work enhances mathematical reasoning in LLMs using an entropy-regularized process reward model (ER-PRM), and we have carefully considered ethical concerns in data usage and algorithmic design. For data sources, our experiments rely on publicly available human preference datasets, ensuring compliance with their respective licenses. These datasets do not involve privacy issues, and we adhere to ethical guidelines in their use. Algorithmically, while ER-PRM improves reasoning stability, reward modeling can influence LLM behavior, potentially reinforcing unintended biases in problem-solving approaches. Additionally, if not carefully managed, preference learning could lead to overfitting on certain reasoning styles, limiting model generalization. To mitigate these risks, we encourage rigorous testing and diverse dataset construction. Lastly, advanced reasoning models pose potential misuse risks, such as academic dishonesty or exploitation in high-stakes decision-making. We recommend ongoing safety evaluations and responsible deployment to ensure robustness and fairness in real-world applications.
ER-PRM has several limitations. High computational cost may hinder scalability, and overfitting risks require careful tuning, particularly in DeepSeek-Math-7B-Instruct. Our approach is mathematics-focused, and its generalization to other domains remains untested. Finally, reward aggregation mechanisms could benefit from more interpretability. Addressing these issues will improve the model’s robustness and applicability.
Appendix B Technical Lemma
@@ Appendix C Best-of-N Evaluation Results in Tables
In this section, we list the performance of the best-of-N evaluation of ER-PRM and other baselines in Tables. The results for the reward models with policy model Mistral-MetaMath-7b are shown in Table 5, which corresponds to Figure 3(a). The results for the reward models with policy model Deepseek-Math-7B-Instruct are shown in Table 6, which corresponds to Figure 3(b).
@@
Appendix D Definition of Step, Partial Chain, and Chain
@@
Appendix E Intuitive Understanding and Theoretical Advantages
E.1 Intuitive Perspective on Entropy-Regularized Process Rewards
We provide an intuitive explanation of entropy-regularized process rewards:
The Soft-Max View: When using initial policy model $\pi_{0}$ (a weaker SFT model like 7B parameter LLM) to generate reasoning paths, our formulation essentially answers the question: "If we continue from this step using our current knowledge, what’s the likelihood of reaching a correct answer?" The entropy regularization term (controlled by $\eta$) determines how optimistic this assessment should be. Higher $\eta$ values make the model focus more on the highest-reward paths, creating a "soft-optimistic" evaluation that weights multiple possible futures according to their potential.
The Soft-Min View: Conversely, when using an optimal policy model $\pi_{*}$ (a strong model like GPT-4o), our formula addresses the question: "What’s the smallest penalty we would need to apply to make this step as good as other alternatives?" This creates a "soft-pessimistic" evaluation that focuses on the worst-case scenarios but with some forgiveness. This approach employs a more cautious strategy to align with the preferences of a stronger model.
Mechanics of Reward Aggregation: To illustrate how our reward aggregation works in practice, consider the step $a^{1}$ in Figure 1. From this step, we see three possible continuation paths reaching $a^{L1,1}$, $a^{L2,2}$, $a^{L3,3}$. We calculate the correctness of each path individually, and get $y_{1}=1$, $y_{2}=1$, and $y_{3}=0$ since the first two paths derive the correct final answer. Once we get the value of different $y_{i}$, we are able to calculate the expected reward score using the equation provided in Figure 1.
E.2 Intuitive Understanding and Theoretical Advantages
Building on this intuitive foundation, we now articulate the key theoretical advantages of our entropy-regularized approach:
(1) Dual Formulation Flexibility:
Our process reward approach offers a significant theoretical advantage through its dual formulation. When sampling completions from the initial policy $\pi_{0}$, our process reward employs a soft-max aggregation as equation equation 2. Conversely, when sampling from the optimal policy $\pi_{*}$, it implements a soft-min aggregation as equation equation 3.
This duality allows our method to adapt smoothly depending on the sampling policy. When sampling from a weaker initial policy $\pi_{0}$, the soft-max formulation acts as an optimistic estimator that amplifies high-reward paths, creating stronger learning signals. When sampling from a stronger optimal policy $\pi_{*}$, the soft-min formulation provides a conservative estimate that focuses on the most reliable paths while maintaining diversity.
(2) Reward Signal Amplification:
Traditional methods calculate rewards as simple expectations under the initial policy $\pi_{0}$: $r(a^{-L})=\mathbb{E}_{\pi_{0}}[r(a|a^{L})]$. While straightforward, this approach can result in weak signals when $\pi_{0}$ has a low probability of sampling high-reward trajectories, which is common in complex mathematical reasoning tasks where correct solutions form a small subset of solution space.
Our entropy-regularized formulation addresses this through the exponential transformation $e^{\eta r(a,x)}$, which amplifies the reward signal before expectation. This transformation effectively redistributes probability mass toward high-reward trajectories, making the learning signal stronger and more informative. The subsequent logarithmic transformation maintains the relative ordering of rewards while controlling the scale.
(3) Independence of Optimal Policy with Single Hyperparameter:
A key theoretical advantage of our method is that it requires tuning only a single hyperparameter $\eta$ that controls the entropy penalty, providing an elegant way to balance the exploitation between high-reward trajectories and the breadth of policy space.
This contrasts with previous approaches that often require complex weighting schemes or multiple parameters to achieve similar effects. Furthermore, our entropy-regularized process reward can be computed using only the initial policy $\pi_{0}$, as demonstrated in equation equation 2. This removes a dependency present in previous RL methods, where the optimal policy is needed.
Appendix F Figure of Experimental Setting
Figure 5: Figurative Illustration of Experimental Setting